A triangle has three vertices: A(3,5), B(7,1), and C(1,-1)
a)Find the equation of the perpendicular bisectors of: AB, BC, and AC
b)Find the coordinates of X, the point of intersection of the perpendicular bisector of AB and the perpendicular bisector of BC

Answer by Alan3354(69443)

A triangle has three vertices: A(3,5), B(7,1), and C(1,-1)
a)Find the equation of the perpendicular bisectors of: AB, BC, and AC
-----
Find the line thru A & B
Slope = diffy/diffx = -4/4 = -1
Slope m of lines perpendicular is the neg inverse, m = 1
---
Find the midpoint: get the average of x & y
x: (3+7)/2 = 5
y: (5+1)/2 = 3
Midpoint (5,3)
Now get the eqn of the line m = 1 thru (5,3)
Use y = mx + b and the point to find b, the y-intercept
3 = 1*5 + b
b = -2
--> eqn is y = x - 2
==========================
Do the same for AC and BC
=======================================
b)Find the coordinates of X, the point of intersection of the perpendicular bisector of AB and the perpendicular bisector of BC
When you have the 2 eqns, solve for the point of intersection.